In the report "Expenditures on Children by Families," the U.S. Department of Agriculture projected the yearly child-rearing expenditures on children from birth through age 17. For a child born in 2010 to a two-parent middle-income family, the report estimated annual expenditures of $10,808 when the child is 6 years old, and $14,570 when the child is 15 years old.
(a) Using the above information, write the two ordered pairs of the form (child's age, annual expenditure). (Let c represent the child's age and y represent annual expenditure.)
(c, y) =
(c, y) =


(b) Assume the relationship between the child's age c and the annual expenditures y is linear. Use your answers to part (a) to write an equation in slope-intercept form that models this relationship.

(c) What are the projected child-rearing expenses when the child is 16 years old?

Answer by ewatrrr(24785)

(15, 14570)
( 6, 10808) m = 3762/9 = 418
.........
y -10808 = 418(x - 6)
y = 418x +8300
.......
x = 16
y = 418(16) +8300 = $14,988